# Calibrated Multi-Preference Optimization for Aligning Diffusion Models

### v2

@@ Abstract
22footnotetext: Work done during an internship at Google DeepMind.33footnotetext: Equal advising.
@@ 1 Introduction
[…]
@@ 4.3 Preference Pair Selection
[…]
@@ 5 Experiments
Table 3: T2I Benchmarks evaluation. We compare the benchmark results of CaPO-SDXL and CaPO-SD3-M on text-to-image benchmarks, e.g., GenEval [16] and T2I-Compbench [21], with various open-source state-of-the-art models (e.g., Flux-dev [29], and Flux-schnell [29], and SE3.5-L [11]). We observe that with CaPO, the majority of evaluation metrics for SDXL and SD3-M show improvement. For comparison, we also include the most recent three image generation models, which are 3$\times$ larger compared to SDXL and SD3-M.
@@ 5.2 Multi-reward experiments
Benchmark results. For quantitative analysis, we evaluate our models on various T2I benchmarks; GenEval [16] which evaluates object-focused generation, and T2I-Compbench [21] for compositional generation. We compare our method with the base models SDXL and SD3-M, as well as open-source state-of-the-art T2I diffusion models such as FLUX-dev [29], FLUX-schnell [29], and Stable Diffusion 3.5-large (SD3.5-L) [11]. Tab. 3 shows that CaPO improves the performance of the base model, e.g., 0.55$\rightarrow$0.59 for SDXL, 0.68$\rightarrow$0.71 on GenEval overall score, and on almost every metrics in T2I-Compbench.
@@ 5.3 Ablation Studies
Table 5: Ablation on loss weighting. We show the results of CaPO multi-reward fine-tuning SDXL with constant weighting (i.e., $-w_{t}\lambda_{t}^{\prime}=1$), and sigmoid weighting by varying bias $b=1.0,1.5,2.0$. Using sigmoid weighting shows better results than constant weighting, and $b=1.5$ performs the best.
@@ A.2 Diffusion preference optimization
¯$\begin{split}\bar{\ell}(\theta)&=g\big{(}R(\mathbf{x}^{+},\mathbf{c})-R(\mathbf{x}^{-},\mathbf{c})\big{)}\\ &=g\bigg{(}\frac{\beta}{2}\mathbb{E}_{t,\boldsymbol{\epsilon}^{+},\boldsymbol{\epsilon}^{-}}\big{[}\lambda_{t}^{\prime}\big{(}\|\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t}^{+};\mathbf{c},t)-\boldsymbol{\epsilon}^{+}\|_{2}^{2}-\|\boldsymbol{\epsilon}_{\textrm{ref}}(\mathbf{x}_{t}^{+};\mathbf{c},t)-\boldsymbol{\epsilon}^{+}\|_{2}^{2}-\|\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t}^{-};\mathbf{c},t)-\boldsymbol{\epsilon}^{-}\|_{2}^{2}+\|\boldsymbol{\epsilon}_{\textrm{ref}}(\mathbf{x}_{t}^{-};\mathbf{c},t)-\boldsymbol{\epsilon}^{-}\|_{2}^{2}\big{)}\big{]}\bigg{)}\\ &\leq\mathbb{E}_{t,\boldsymbol{\epsilon}^{+},\boldsymbol{\epsilon}^{-}}\bigg{[}g\bigg{(}\tfrac{1}{2}\beta\lambda_{t}^{\prime}\big{(}\|\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t}^{+};\mathbf{c},t)-\boldsymbol{\epsilon}^{+}\|_{2}^{2}-\|\boldsymbol{\epsilon}_{\textrm{ref}}(\mathbf{x}_{t}^{+};\mathbf{c},t)-\boldsymbol{\epsilon}^{+}\|_{2}^{2}-\|\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t}^{-};\mathbf{c},t)-\boldsymbol{\epsilon}^{-}\|_{2}^{2}+\|\boldsymbol{\epsilon}_{\textrm{ref}}(\mathbf{x}_{t}^{-};\mathbf{c},t)-\boldsymbol{\epsilon}^{-}\|_{2}^{2}\big{)}\bigg{)}\bigg{]}\text{,}\end{split}$
@@ A.3 Loss weighting
Figure 6: Training dataset score distribution. We plot the histogram of rewards (top row) and calibrated rewards (bottom row) of training dataset. By using calibration, the scores are centered and bounded in range [0,1][0,1][ 0 , 1 ].
@@ B.2 Training and evaluation
[…]
@@ Appendix C Additional ablation study
Loss weighting for SD3-M. We show the effect of loss weighting when training SD3-M models. Similar to SDXL, we compare the results of CaPO multi-reward fine-tuning with different bias parameters. In Tab. 6, we show that sigmoid weighting with bias $b=-1.0$ shows the best result, outperforming the constant weighting counterpart. Note that for SDXL, $b=1.5$ performs the best, while for SD3-M, negative bias $b=-1.0$ performs the best. Remark that as SD3-M performs diffusion modeling on 16×128×12816$\times$ 128$\times$ 12816 × 128 × 128, and SDXL performs on 4×128×1284$\times$ 128$\times$ 1284 × 128 × 128, the bias shifts toward negative as the total variance becomes higher, and the log-SNR should be increased [18, 19].
@@
Sigmoid weighting ($b=-1.0$) 59.0 55.7 69.3
Table 6: Ablation on loss weighting for SD3-M. We show the results of CaPO multi-reward fine-tuning SD3-M with constant weighting (i.e., $-w_{t}\lambda_{t}^{\prime}=1$), and sigmoid weighting by varying bias $b=0.0,-1.0$. Similar to SDXL, using sigmoid weighting shows better results than constant weighting, and $b=-1.0$ performs the best.

### v3

@@ Abstract
22footnotetext: Work done during an internship at Google DeepMind.footnotetext: Corresponding Authors.  $\ddagger$ Equal advising.
@@ 1 Introduction
Figure 2: Overview. (a) We generate $N$ images using pretrained T2I diffusion model using the prompt dataset, and infer the scores from reward models. (b) Then, we calibrate the rewards by making pairwise comparison between images. For each image, we compute the win-rates between other N1N-1 images using Eq. (2), and average them to obtain calibrated reward $R_{\textrm{ca}}$ (see Sec. 4.2). (c) We select pair by choosing the best-of-NN and worst-of-NN when using single reward. For multi-reward, we use non-dominated sorting algorithm to select upper Pareto set as positives, and lower Pareto set as negatives. The accepted and rejected pairs are also listed using proposed rejection sampling method. (d) Lastly, during training, we select a pair from (c), and compute CaPO loss (i.e., Eq. (8)), which perform regression task to match the difference in calibrated rewards, i.e., $\Delta R_{\textrm{ca}}$ by the difference of implicit reward model $\Delta R_{\theta}$.
@@ 4.3 Preference Pair Selection
The best-of-$N$ sampling [42, 8] or rejection sampling [62] methods that select samples with highest reward from $N$ generation are commonly used in RLHF. For a single reward, it is straightforward to choose the sample $\mathbf{x}^{+}$ with highest reward, and $\mathbf{x}^{-}$ that has lowest reward to maximize the margin between the pair. For multi-reward optimization, the naïve approach is to use weighted sum as the total proxy reward model, and perform rejection sampling with it. However, choosing the weights often relies on heuristics, and the optimal weights might be dynamic depending on the input, which can lead to suboptimal performance.
@@ 5 Experiments
[…]
@@ 5.2 Multi-reward experiments
[…]
@@ 5.3 Ablation Studies
Table 5: Ablation on loss weighting. We show the results of CaPO multi-reward fine-tuning SDXL with constant weighting (i.e., $w_{t}\lambda_{t}^{\prime}=1$), and sigmoid weighting by varying bias $b=1.0,1.5,2.0$. Using sigmoid weighting shows better results than constant weighting, and $b=1.5$ performs the best.
@@ A.2 Diffusion preference optimization
$\begin{split}\bar{\ell}(\theta)&=g\big(R(\mathbf{x}^{+},\mathbf{c})-R(\mathbf{x}^{-},\mathbf{c})\big)\\ &=g\bigg(\frac{\beta}{2}\mathbb{E}_{t,\boldsymbol{\epsilon}^{+},\boldsymbol{\epsilon}^{-}}\big[\lambda_{t}^{\prime}\big(\|\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t}^{+};\mathbf{c},t)-\boldsymbol{\epsilon}^{+}\|_{2}^{2}-\|\boldsymbol{\epsilon}_{\textrm{ref}}(\mathbf{x}_{t}^{+};\mathbf{c},t)-\boldsymbol{\epsilon}^{+}\|_{2}^{2}-\|\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t}^{-};\mathbf{c},t)-\boldsymbol{\epsilon}^{-}\|_{2}^{2}+\|\boldsymbol{\epsilon}_{\textrm{ref}}(\mathbf{x}_{t}^{-};\mathbf{c},t)-\boldsymbol{\epsilon}^{-}\|_{2}^{2}\big)\big]\bigg)\\ &\leq\mathbb{E}_{t,\boldsymbol{\epsilon}^{+},\boldsymbol{\epsilon}^{-}}\bigg[g\bigg(\tfrac{1}{2}\beta\lambda_{t}^{\prime}\big(\|\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t}^{+};\mathbf{c},t)-\boldsymbol{\epsilon}^{+}\|_{2}^{2}-\|\boldsymbol{\epsilon}_{\textrm{ref}}(\mathbf{x}_{t}^{+};\mathbf{c},t)-\boldsymbol{\epsilon}^{+}\|_{2}^{2}-\|\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t}^{-};\mathbf{c},t)-\boldsymbol{\epsilon}^{-}\|_{2}^{2}+\|\boldsymbol{\epsilon}_{\textrm{ref}}(\mathbf{x}_{t}^{-};\mathbf{c},t)-\boldsymbol{\epsilon}^{-}\|_{2}^{2}\big)\bigg)\bigg]\text{,}\end{split}$
@@ A.3 Loss weighting
Figure 6: Training dataset score distribution. We plot the histogram of rewards (top row) and calibrated rewards (bottom row) of training dataset. By using calibration, the scores are centered and bounded in range $[0,1]$.
@@ B.2 Training and evaluation
Training configuration. Throughout experiments, we use Jax [14] and train models using the Optax library on TPU chips. For both SDXL and SD3-M experiments, we use Adam [26] optimizer. Regarding training configuration for SDXL experiments, we use batch size of 256, learning rate of 1e-5 with linear warmup for first 1000 steps, and train for maximum 10000 steps. For SD3-M, we use batch size of 256, learning rate of 1.5e-5 with linear warmup for first 1000steps, and train for maximum 5000 steps. We choose hyperparameter $\beta$ by sweeping over $\{300,500,1000\}$ for CaPO, $\{500,1000,2000\}$ for IPO, and $\{2000,3000,4000\}$ for DPO when training SDXL model. For SD3-M, we sweep over $\{30,50,100\}$ for CaPO, $\{50,100,200\}$ for IPO, and $\{100,200,300\}$ for DPO. For all training, we use sigmoid loss weighting with $b=1.5$ for SDXL and b=1.0b=-1.0 for SD3-M (including all DPO, IPO, and CaPO). During training, we generate images using subset of Parti prompts at each 1000-th iteration, and choose the final model with maximum validation win-rate (average of win-rates for multi-reward signals).
@@ Appendix C Additional ablation study
Loss weighting for SD3-M. We show the effect of loss weighting when training SD3-M models. Similar to SDXL, we compare the results of CaPO multi-reward fine-tuning with different bias parameters. In Tab. 6, we show that sigmoid weighting with bias b=1.0b=-1.0 shows the best result, outperforming the constant weighting counterpart. Note that for SDXL, $b=1.5$ performs the best, while for SD3-M, negative bias b=1.0b=-1.0 performs the best. Remark that as SD3-M performs diffusion modeling on $16\times 128\times 128$, and SDXL performs on $4\times 128\times 128$, the bias shifts toward negative as the total variance becomes higher, and the log-SNR should be increased [18, 19].
@@
Sigmoid weighting (b=1.0b=-1.0) 59.0 55.7 69.3
Table 6: Ablation on loss weighting for SD3-M. We show the results of CaPO multi-reward fine-tuning SD3-M with constant weighting (i.e., $w_{t}\lambda_{t}^{\prime}=1$), and sigmoid weighting by varying bias b=0.0,1.0b=0.0,-1.0. Similar to SDXL, using sigmoid weighting shows better results than constant weighting, and b=1.0b=-1.0 performs the best.